how do i set up the problem :
Audrey has 32 dimes and quarters in her piggy bank with a total value of 6.50. How many of the coins are dimes and how many are quarters?
I need to write 2 equations and solve by addition method

Found 2 solutions by ewatrrr, nerdybill:
Answer by ewatrrr(24785)


Hi,
I need to write 2 equations and solve by addition method
in that case, let x = number of dimes and y = number of quarters
x + y = 32
10x + 25y = 650 cents
Multiplying both sides of the first equation by -25 to eliminate y
-25x -25y = 32*(-25)
10x + 25y = 650 cents
-15x = -150
x = 10, the enumber of dimes
the quarters would be 22 (32-10)
checking our answer
10*10 + 22*25 = 100 + 550 = 650cents Or $6.50

Answer by nerdybill(7384)

Audrey has 32 dimes and quarters in her piggy bank with a total value of 6.50. How many of the coins are dimes and how many are quarters?
I need to write 2 equations and solve by addition method
.
Let x = number of dimes
and y = number of quarters
.
From: "Audrey has 32 dimes and quarters in her piggy bank"
x + y = 32 (equation 1)
.
From: "total value of 6.50"
.10x + .25y = 6.50 (equation 2)
.
Using the "addition method", to eliminate x, we multiply both sides of equation 1 by -.10 and "add" the two equations together:
-.10x - .10y = -3.2
.10x + .25y = 6.50
----------------------
.15y = 3.3
y = 22 (number of quarters)
.
Substitute the above value of y back into equation 1 to find x:
x + y = 32
x + 22 = 32
x = 32-22
x = 10 (number of dimes)